Solve for X. Simplify your answer use comma to separate answer as needed.
4x(x-9)-5x(x-8)= -5

Answer by algebrahouse.com(1659)

4x(x - 9) - 5x(x - 8) = -5
4x^2 - 36x - 5x^2 + 40x = -5 {used distributive property}
-x^2 + 4x = -5 {combined like terms}
x^2 - 4x - 5 = 0 {added x^2 and subtracted 4x from both sides}
(x - 5)(x + 1) = 0 {factored into two binomials}
x - 5 = 0 or x + 1 = 0 {set each factor equal to 0}
x = 5 or x = -1 {solved each equation for x}
x = 5,-1
www.algebrahouse.com